If a random variable u has t -distribution with n degree of freedom, find the 

distribution of u^2.


1
Expert's answer
2021-05-07T11:43:28-0400

Let s shows the sample standard deviation, "\mu" shows the populaiton mean, "\bar{x}" shows the sample mean, "\sigma" is population standard deviation. and n+1 is sample size.


Chi sqaure distribution with (n-1) degree of freedom will be


"\chi^{2}_{(n-1)}\sim \frac{s^{2}(n-1)}{\sigma^{2}}"


Here we will use following relationships between the variables:


Relation between standard normal distribution Z and chi-sqaure distribution:


"Z^{2}\sim \chi^{2}_{1}"


Relation between F distribution and chi-sqaure distribution:


"F_{a,b}\equiv \frac{\chi^{2}_{a}/a}{\chi^{2}_{b}/b}"


The t-statistics with n degree of freedom will be


"t_{n}=\frac{\bar{x}-\mu}{s/\sqrt{n+1}}=\frac{\bar{x}-\mu}{\sigma/\sqrt{n+1}}\cdot \frac{\sigma}{s}=Z\cdot \frac{\sigma}{s}=\frac{Z}{\sqrt{\frac{s^{2}}{\sigma^{2}}}}=\frac{Z}{\sqrt{\frac{s^{2}(n)}{\sigma^{2}(n)}}}=\frac{Z}{\sqrt{\frac{\chi^{2}_{n}}{n}}}"


 "=\frac{\sqrt{\frac{\chi^{2}_{1}}{1}}}{\sqrt{\frac{\chi^{2}_{n}}{n}}}"


Now squaring both sides gives:


"U=t^{2}_{n}=\frac{\frac{\chi^{2}_{1}}{1}}{\frac{\chi^{2}_{n}}{n}}\sim F_{1,n}"


Hence, "u^2" has F distribution.